OpenStudy (anonymous):

Find the height of a square pyramid that has a volume of 12 cubic feet and a base length of 3 feet. (1 point) Select one: a. 1.5 feet b. 2 feet c. 3 feet d. 4 feet

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

V=1/3*b^2*h 12=1/3*3^2*h 12=1/3*9*h 12=3*h h=12/3 h=4 feet